Summary of the Meeting of
the PS Management Board
held on 22.11.2002
Present: J. Boillot, M. Chanel, G. Daems,
J.P. Delahaye (Chairman), B. Frammery, R. Garoby, G. Geschonke, J. Gruber,
H. Haseroth, T. Kehrer (Secretary), J. Lettry, E. Métral (Staff Association),
K.D. Metzmacher, J.P. Potier
Excused: M. Lindroos, S.L. Neboux, K. Schindl
* * * * *
1. Follow-up of previous
meeting and matters arising
K. Schindl will report in the next PSMB on the workshop that took place
on 7-8.11.2002 at CERN on the LIS Project. A written summary
of the meeting was sent to the PSMB.
Studies (0.6 sec repetition time study
on the Booster): this subject will be discussed in today's PPC
meeting and M. Benedikt will report on the study in the next PSMB.
PS for AB-CO Group: J.P. Delahaye congratulated
the CO group on the organization of the visit of the PS installations
last Wednesday. All the members of the future AB-CO group were invited.
The success of the visit was also due to the excellent guides.
Extension of operation: the request to
extend the operation of CTF2 from 29.11.2002 until 13.12.2002 was accepted.
- Shutdown work: it was felt that the schedules of all shutdown
work/interventions (including tests on physics experiments) should be
distributed and/or published in order to facilitate the coordination
in the different service groups.
for AB Technical Meetings. J.P. Delahaye
presented the proposed structure. Group Leaders are encouraged to give
their comments and suggestions.
Publications: In spite of a very successfull operation in PS division
over the last ten years, the ABMB is not in favour of an Editorial Board
and it was agreed that the Group/Project leaders should approve the
papers before submission to the Division Leader. Abstracts of papers
to be submitted to big Accelerator Conference will be reviewed at the
ABMB. Preparation of PAC03 papers will follow the actual PS procedure.
SL and PS Staff Members will receive a letter announcing their transfer
to the AB Division as from 1.1.2003.
- PS Parties: there will be no PS Party this
year. Instead each group is allowed to spend up to 15 CHF/pp for the
organization of internal events. The reimbursement of these expenses
will be processed by the Divisional Secretariat upon presentation of
receipts. The amounts per group will be communicated by J. Boillot (action
Control Room: S. Baird expressed his concerns
about the advancement of the study for the future control room. Nothing
new since the last meeting that took place in the beginning of October
(S. Baird, R. Bailey, C. Wyss and the architect). S. Baird contacted
C. Wyss on several occasions about this subject, without success. J.P.
Delahaye reported from the last ACMB that the negotiations to obtain
the wooden sphere are still going on and he will contact C. Wyss.
submission of abstracts to the Editorial Board: 26.11.2002.
Info: suggestions for subjects to
Eveline Durieu-Thiry: 01.12.2002.
Report: contributions to B.W. Allardyce:
2. Shutdown work
tests: K. Metzmacher pointed out, that tests with the new thyristor
gate drives and regulation electronics for the PFWs were performed
during two weeks after the physics run. At high power levels spurious triggers
happened, problems that could not be resolved. Further tests with a high
power test load are required and therefore the existing electronics will
remain for the 2003 run.
31: the new septum/tank SES 31 assembly was recently tested in the lab.
During final inspection, breakdown traces on a ceramic insulator, opposite
an inner tank weld, were found. The tank details were reviewed, the design
adapted and the tank modified. The septum assembly and testing procedure
will have to be repeated. If not ready in time, the old SES 31 will remain
in the PS in 2003.
57 : in the PS, SMH 57 was routinely tested for leaks. A leak from the
hydraulic circuit into the vacuum was found (for this, the cooling water
is replaced by helium at the same pressure, i.e. 15 bars). Its operational
spare will replace septum SMH 57.
3. Report from the SPSC Meeting held on 15.11.02 (J.P.
Riunaud) and the Research Board meeting held on 12.11.02 (J.P. Delahaye).
J.P. Riunaud gave a summary
of the 60th SPSC meeting held on 15.11.02, which included status reports
on SPS, PS and PS, SPS experiments, followed by status reports on experiments.
to the ISOLDE team! It was an excellent year with the parallel operation
of both the GPS and the HRS for the first time and a record for the
number of shifts (375).
: no extension for DIRAC for the moment. This situation can be reviewed
in February if supported by the next SPSC (possible conflict with n-TOF
since DIRAC and n-TOF cannon use the same PS cycle).
project to test the use of antiprotons for medical treatment is under
study and will be presented at the next SPSC. The Project may not fulfil
the required CERN physics criteria.
2003 was adopted.
42 has been approved. The CERN Management Board thanked H.Haseroth
for all the work he did as a Chairman for SAPOCO. H. Haseroth will continue
this duty next year.
of the December Committees:
of the LTP (Long Term Plan).
Recruited Staff (LS): pending approval by the Council. This recruitment
will be subject to Selection Boards as for any other vacancy and
will be based on the actual career paths A and B. A change to career
path C could be envisaged later on. Conditions of this career change
as well as other contractual matters related to the status of LS
are still to be approved in the next Finance Committee.
4. Status of the PS budget (J. Boillot)
J. Boillot presented the status
of the PS budget on 20 November 2002.
- The commitments now reach 100% and the payments exceed
60% of the total budget (Exploitation and Projects). Provisions will be
decided in the bookclosing exercise.
- 120 KCHF were credited on the CLIC/CTF budget in November
(from klystron repair insurance)
- 440 KCHF from a loan to ASACUSA in previous years will
be credited next year (340 KCHF on CLIC/CTF and 100 KCHF on other budget
- The CLIC budget will be fully committee after increasing
the present commitment by 1 MCHF which corresponds to one third of the cost
of the accelerating cavities (contract signed in 2002).
Following a request by S. Myers to FI Division, it should
be possible to finance three posts for three years with budget saved from
the material budget in 2002 and transferred in 2004, 2005 and 2006 to the
Personnel budget. Priority is given to PO, RF and BT.
5. Jobs of the Operators during the shutdown (S. Baird)
S. Baird presented the list of operators
and the work that has been assigned to them during the shutdown. The list
will be sent to the Group Leaders and any additional request should be sent
to S. Baird. Once the list finalized, a memo will inform operators and supervisors
of their duties. The supervisor will have to participate in the MAPS exercise
of his supervisee.
6. PS Consolidation Project (J. Gruber)
Due to lack of time, it was not possible
to discuss this subject at the meeting. J. Gruber will send his Draft report
to the PSMB for approval.
Division | PSMB
Meetings | Other
PS Meetings |
22 November, 2002